John mixed cashews and almonds. John bought 4 pounds of almonds for a total cost of $22. The cost per pound of cashews is 60% more than the cost per pound for almonds. John bought enough cashews that, when he mixed them with the almonds, the mixture had a value of $6.50 per pound. Aprox what percent of the mixture, by weight, was cashews?
I figured the cost per pound for almonds is $5.50 and cost per pound for cashews is $8.80 but having problems with mixture question.

Answer by rothauserc(4718)

4 * 5.50 = 22
x * 8.80 = 8.80x
(4+x) * 6.50 = 6.50(4+x)
therefore we have
22 + 8.80x = 6.50(4+x)
22 + 8.80x = 26 + 6.50x
2.30x = 4
x = 1.739130435 approx 1.74 lbs of cashews
percent of the mixture, by weight, for cashews =
1.739130435 / (4 + 1.739130435) = 0.303030303 approx 30%